OpenStudy (pgreene008):

Analyze the function f(x) = sec 2x. Include: - Domain and range - Period and Amplitude - Two Vertical Asymptotes

OpenStudy (pgreene008):

@nincompoop

OpenStudy (pgreene008):

@Nnesha

OpenStudy (cathyangs):

What are those things supposed to be 'included' in? A graph or a verbal explanation?

OpenStudy (pgreene008):

like written down

OpenStudy (cathyangs):

Okay. Knowing what sec is (1/cos, in case you forgot), what is the possible domain, just around 0, for this function? if cos(x) is equal to 0, there's no value for the function sec(x) because you can't divide by 0. That's where asymptotes form.

OpenStudy (pgreene008):

so the domain would be: all real numbers except π/4 + πn ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but close cosine is zero at \(\frac{\pi}{2}\) not \(\frac{\pi}{4}\)

OpenStudy (cathyangs):

If you want, @satellite73 , you can take over.

OpenStudy (pgreene008):

so what ould it be @satellite73 and thanks for your help @cathyangs

OpenStudy (cathyangs):

Or not? @pgreene008 you may be stuck with me XD. So we've got the domain, and thus also know where asymptotes form. The period asks the length of the x-axis segments before the function repeats. You can find this by graphing or by using the equation breakdown of this struture ( https://mathway.com/examples/Trigonometry/Graphing-Trigonometric-Functions/Amplitude-Period-and-Phase-Shift?id=342) whichever style works for you. Finally, the amplitude can be found either way above, as well. Let me know if this all makes sense!

OpenStudy (alekos):

If you use Desmos you'll get all the info you require

OpenStudy (pgreene008):

so would this be the domain:?? Domain: all real numbers other than: ±π/4, ±3π/4, ±5π/4 @alekos @cathyangs

OpenStudy (alekos):

Yes, that's the domain

OpenStudy (alekos):

and so on i.e. ±(2n-1)π/4 where n = 1,2,3..........

OpenStudy (pgreene008):

so what would the final answer be?

OpenStudy (alekos):

Domain is all real numbers except x = ±(2n-1)π/4 where n = 1,2,3..........

OpenStudy (alekos):

The function is undefined at these points

OpenStudy (anonymous):

A plot is attached.